OpenStudy (anonymous):

IN 5 YEARS JOHN WILL BE 25YEAR OLDER THAN CHARLES WILL BE THEN 2YEARS AGO JOHN WAS SEVEN TIMES AS OLD AS CHARLES WAS LAST YEAR. HOW OLD ARE THEY?

OpenStudy (mertsj):

j+5 = john in 5 years c+5 = charlie in 5 years j+5 = c+5 +25 j-2 = john 2 years ago c-1 = charlie last year j-2=7(c-1)

OpenStudy (mertsj):

Now solve the two equations simultaneously

OpenStudy (anonymous):

THANKS BUT SOLVE IT FOR ME

OpenStudy (mertsj):

Are you sure you have the problem down exactly as written? I am getting a fraction and usually these problems turn out to be whole numbers?

OpenStudy (anonymous):

Er.. I think there is an error in the question

OpenStudy (anonymous):

I get Charles -> 5 1/3 John -> 32 1/3

OpenStudy (mertsj):

\[j+5=c+30\]

OpenStudy (anonymous):

typo?

OpenStudy (mertsj):

\[j-c=25\]

OpenStudy (mertsj):

\[j-2=7(c-1)\]

OpenStudy (mertsj):

\[j=7c-5\]

OpenStudy (mertsj):

\[7c-5+5=c+30\]

OpenStudy (mertsj):

6c=30 c=5 j=7(5)-5 j=30